OpenStudy (anonymous):

Find a possible solution to the equation sin(3x+13)=cos(4x). A. x=-13 degrees B. x=0 degrees C. x=11 degrees D. x-13 degrees

OpenStudy (welshfella):

just replace x in the equation by the values of x in the options and see which one fots

OpenStudy (welshfella):

* fits

OpenStudy (anonymous):

hint: cos(4x) = sin(90 - 4x)

OpenStudy (anonymous):

still dont get it

OpenStudy (anonymous):

sin(3x+13)=cos(4x) sin(3x+13)=sin(90 - 4x) this implies 3x + 13 = 90 - 4x solve for x

OpenStudy (anonymous):

x=11

OpenStudy (welshfella):

right